OpenStudy (anonymous):

Use the Pythagorean Theorem to determine the distance between points A(5, 3) and B(-3, -3).

OpenStudy (anonymous):

heaven forbid we should call it the 'distance formula'!

OpenStudy (anonymous):

D = sqrt ( (D in x)^2 + (D in y)^2) D = sqrt ( (5 - (-3))^2 + (3 - (-3))^2 ) D = sqrt ( (8)^2 + (6)^2) D = sqrt (64 + 36) D = 10